# Monotonic Path-Specific Effects: Application to Estimating Educational Returns

### v1

@@ 1 Introduction
One of the most resilient social scientific findings across a range of national contexts is the strong association between educational attainment and a variety of life outcomes, including earnings, health, social capital, and family stability (Hout, 2012; Chetty et al., 2023). Conventionally, researchers have taken one of two approaches to evaluating the social and economic returns to education: the first employs a “years of schooling” measure of educational attainment (Angrist and Krueger, 1991, 1992; Kane and Rouse, 1993; Card, 1994; Ashenfelter and Zimmerman, 1997; Card, 1999, 2001; Angrist and Chen, 2011), while the second dichotomizes attainment as a point-in-time treatment. This latter approach has been particularly influential in the study of the impact of postsecondary attainment on earnings, where the treatment considered is often an indicator for whether an individual has attended, or graduated from, college (Brand and Xie, 2010; Carneiro et al., 2011; Zimmerman, 2014; Goodman et al., 2017; Smith et al., 2020; Bleemer, 2022; Mountjoy, 2022).
@@
Despite the important insights this literature has made into establishing the causal effect of educational attainment on important social and economic outcomes, extant work has been inattentive to the sequential process by which people make educational transitions (Mare, 1980). [^1] I use the term “educational transition” to refer both to vertical transitions (e.g. enrollment at a secondary or tertiary institution), as well as to the attainment of a qualification at a given level (e.g. high school graduation or BA completion). At the end of high school, individuals decide whether or not to enroll in college. Among college enrollees, only 60% receive a BA within six years of initial college entry (Snyder et al., 2016), with an even lower proportion for low-income students and students of color (Eller and DiPrete, 2018; Zhou and Pan, 2023). Moreover, amidst higher educational expansion in the US, college graduates must increasingly choose whether to enter the labor market or to enroll in post-graduate education. Increasingly, therefore, educational attainment in the US has become a field of multiple levels with sequential transitions, all of which are independently consequential for individuals’ labor market outcomes, and therefore of independent scientific interest.
@@
The sequential nature of educational transitions implies that a causal mediation framework can be employed to study the causal paths by which education’s “value-added” occurs. Specifically, we can consider the first transition in a sequence of educational levels of interest as a treatment variable, $A$, and subsequent transitions as mediators that “transmit” the effects of the treatment and of prior transitions, $M_{k}$ ($1\leq k\leq K$). For example, if we are interested in the total effect of high school completion on earnings, we may ask to what extent this total effect operates indirectly, through the effects of college attendance and college completion (putative mediators) on earnings, or directly, through alternative causal pathways. The insight that the total causal effect of education can be decomposed into its direct and indirect effects opens up a range of important research and policy-oriented questions. For example, tracing to what extent an early-stage educational intervention boosts outcomes such as earnings via its promotion of subsequent educational attainment (its indirect effects), or via earnings directly, would enable policy-makers to discern what drives the intervention’s value and to hone subsequent policy (e.g. Hurwitz and Howell, 2014; Sullivan et al., 2019; Castleman et al., 2020; Bird et al., 2021; Dynarski et al., 2021; Black et al., 2023; Turner and Gurantz, 2024). Relatedly, if the early intervention’s effects are heterogeneous across demographic groups, assessing the intervention’s direct and indirect effects could guide researchers to aspects of the educational experience that either promote or inhibit upward mobility. Nevertheless, prior empirical approaches are not well-suited to answering these questions: a “years-of-schooling” approach captures the direct effect of each additional year of schooling, while the dichotomous approach conflates the direct and indirect effects. [^2] A further strand of literature, especially prominent in labor economics, explores labor market returns to horizontal aspects of differentiation within a given educational level (e.g. college selectivity, as well as specific colleges) or college types (e.g. Cohodes and Goodman, 2014; Goodman et al., 2017; Mountjoy and Hickman, 2021; Chetty et al., 2023; Eller, 2023). While my proposed framework prioritizes the effects of different levels of education, I discuss in my concluding remarks how the framework could be extended to accommodate multivariate mediators.
@@ 2.1 A Single Intermediate Transition
Using potential outcomes notation, let $M(a)$ denote an individual’s potential value of the mediator if their treatment status were set to $a$, and let $Y(a,m)$ denote that individual’s potential outcome if their treatment and mediator statuses were set to $a$ and $m$, respectively. I assume, as I formalize in the following section, that sequential transitions are characterized by monotonicity; here, this means that individuals who do not complete high school cannot attend college, or $M(0)=0$. This sequential nature of educational transitions therefore implies the following set potential outcomes: $\{Y(1),Y(0),Y(1,0),Y(1,1)\}$. Further, since by the composition assumption $Y(a)=Y(a,M(a))$ (VanderWeele and Vansteelandt, 2009), under monotonicity we have that Y(0)=Y(0,M(0))Y(0)=Y(0,M(0))=Y(0,0)=Y(0,0), and that
$Y(1)$ $=Y(1,0)+M(1)[Y(1,1)-Y(1,0)].$
Thus, the individual total effect of $A$ on $Y$ can be decomposed as
$Y(1)-Y(0)$ $=Y(1,0)-Y(0,0)+M(1)[Y(1,1)-Y(1,0)].$ (1)
Since these individual-level quantities are unidentified, I focus on their population-level analogs. Taking the expectation of Equation 1, we obtain the following decomposition of the ATE (see Zhou, 2022a):
@@
$=\underbrace{\Delta_{0}}_{A\to Y}+\underbrace{\pi_{\textup{1}}\Delta_{\textup{1}}+\eta_{\textup{1}}}_{A\to M\to Y}.$ (3)
@@
Here, $\Delta_{\textup{0}}$ and $\Delta_{1}$ denote the direct effects of the first and intermediate transitions on the outcome, $A\rightarrow Y$ and $M\rightarrow Y$, respectively, $\pi_{\textup{1}}$ denotes the total effect of the first transition on the intermediate transition $A\to M$, and $\eta_{\textup{1}}$ denotes the covariance between the effect of the initial transition on completion of the second and the effect of the second transition on $Y$. Specifically, $\eta_{\textup{1}}$ is positive if those who would attend college given high school completion (i.e., $M(1)=1$) benefit more from college attendance in terms of their later earnings (i.e., have a larger Y(1,1)Y(1,0)Y(1,1)-Y(1,0)) than those who do not (i.e., $M(1)=0$), and negative if the opposite is true. Meanwhile, the composite term $(\pi_{\textup{1}}\Delta_{1}+\eta_{\textup{1}})$ captures the average indirect effect of the treatment via the intermediate transition ($A\rightarrow M\rightarrow Y$), comprising the sum of (i) the probability of college enrollment if an individual graduated high school, multiplied by the direct of college enrollment, and (ii) the covariance between college enrollment and its direct effect on earnings.
@@ 2.2 Generalization to $K$ Intermediate Transitions
I now generalize the approach introduced in the preceding section to the case of $K$ intermediate transitions. As previously, I denote the treatment (“initial transition”) of high school graduation by $A$, and use $M_{1},\dots M_{K}$ to refer to the $K$ subsequent transitions of interest (“intermediate transitions”), where I assume that all of $M_{1},\dots,M_{K}$ are binary and that for any $i<j$, $M_{i}$ temporally precedes $M_{j}$. For instance, we may wish to decompose the total effect of high school completion on earnings via college attendance ($M_{1}$), college completion ($M_{2}$) and graduate school attendance ($M_{3}$). Let an overbar denote a vector of variables, such that $\overline{M}_{k}=(M_{1},M_{2},\dots M_{k})$ and $\overline{1}_{k}=(A=1,M_{1}=1,\dots,M_{k-1}=1)$. Further, let $[K]$ denote the set $\{0,1,\dots,K\}$. In addition, I denote by $X$ a vector of pretreatment confounders of the effect of $(A,\overline{M}_{k})$ on ($M_{k+1},Y)$, and by $\overline{Z}_{k}=(Z_{1},\dots Z_{k})$ a vector of intermediate confounders that may confound the causal effect of $M_{k}$ on ($M_{k+1},Y)$. Using potential outcomes notation, $Y(\overline{1}_{k},m_{k})$ thus denotes an individual’s potential earnings if they completed, possibly contrary to fact, the treatment in addition to k1k-1 intermediate transitions, and then either completed ($m_{k}=1)$ or did not complete ($m_{k}=0$) the $k$th intermediate transition. Similarly, $M_{k+1}(\overline{1}_{k+1})$ denotes an individual’s potential value of the $k+1$th intermediate transition were that individual to complete the treatment as well as $k$ prior intermediate transitions. As is standard in the mediation literature, I make the following composition assumption (VanderWeele and Vansteelandt, 2009):
@@
Composition: $Y(\overline{1}_{k},m_{k})=Y(\overline{1}_{k},m_{k},M_{k+1}(\overline{1}_{k},m_{k})),\forall k\in[K-1],M_{0}\equiv A$.
In words, Assumption 1 states that a person’s potential outcome under $(\overline{1}_{k},m_{k})$ is equal to their potential outcome under $A=1,\dots,M_{k-1}=1,m_{k}$ and under the value $M_{k+1}$ would naturally take under $A=1,\dots,M_{k-1}=1,m_{k}$. I also invoke the following constraint on units’ potential transition values:
@@
Informally, Assumption 2 (monotonicity) states that an individual’s potential $k+1$th transition value is deterministically 0 if that individual fails to complete the prior ($k$th) transition. It is analogous to a one-sided non-compliance assumption within an instrumental variables (IV) framework, which precludes the presence of both “defiers” as well as “always-takers” principal strata. We can then use this assumption to decompose the ATE of $A$ on $Y$, which I denote by $\tau_{0}$. Specifically, let $\tau_{k}$ denote the gross effect of the $k$th mediator on $Y$, i.e.,
$\tau_{k}=\mathbb{E}[Y(\overline{1}_{k+1})-Y(\overline{1}_{k},0)],$
@@
$\Delta_{k}=\mathbb{E}[Y(\overline{1}_{k+1},0)-Y(\overline{1}_{k},0)].$
To explicate my approach, note that the gross effect of the $k$th mediator, $\tau_{k}$, includes not only the direct effect $M_{k}\rightarrow Y$, net of subsequent educational transitions $\Delta_{k}$, but also the indirect effects of $M_{k}$ via subsequent transitions ($M\rightsquigarrow Y$, where a squiggly arrow denotes a combination of multiple paths). This insight motivates us to further decompose $\tau$ into its direct and indirect components. Under the composition assumption, $\tau_{k}$ can be decomposed as
@@
$\pi_{k+1}$ $=\mathbb{E}[M_{k+1}(\overline{1}_{k+1})],$
$\eta_{k+1}$ $=\textup{cov}[M_{k+1}(\overline{1}_{k+1}),Y(\overline{1}_{k+2})-Y(\overline{1}_{k+1},0)].$

### v2

@@ 1 Introduction
One of the most resilient social scientific findings across a range of national contexts is the strong association between educational attainment and a variety of life outcomes, including earnings, health, social capital, and family stability (Hout, 2012; Chetty et al., 2023). Conventionally, researchers have taken one of two approaches to evaluating the social and economic returns to education: the first employs a “years of schooling” measure of educational attainment (Angrist and Krueger, 1991, 1992; Kane and Rouse, 1993; Card, 1994; Ashenfelter and Zimmerman, 1997; Card, 1999, 2001; Angrist and Chen, 2011), while the second dichotomizes attainment as a point-in-time treatment. This latter approach has been especially influential in the study of the impact of postsecondary attainment on earnings, where the treatment considered is often an indicator for whether an individual has attended, or graduated from, college (Brand and Xie, 2010; Carneiro et al., 2011; Zimmerman, 2014; Goodman et al., 2017; Smith et al., 2020; Bleemer, 2022; Mountjoy, 2022).
@@
Despite the important insights this literature has made into establishing the causal effect of educational attainment on important social and economic outcomes, extant work has been inattentive to the sequential process by which people make educational transitions (Mare, 1980). [^1] I use the term “educational transition” to refer both to vertical transitions (e.g. enrollment at a secondary or tertiary institution), as well as to the attainment of a qualification at a given level (e.g. high school graduation or BA completion). At the end of high school, individuals decide whether or not to enroll in college. Among college enrollees, only 60% receive a BA within six years of initial college entry (Snyder et al., 2016), with an even lower proportion for low-income students and students of color (Eller and DiPrete, 2018; Zhou and Pan, 2023). Moreover, amidst higher educational expansion in the US, college graduates must increasingly choose whether to enter the labor market or to enroll in postgraduate education. Increasingly, therefore, educational attainment in the US has become a field of multiple levels with sequential transitions, all of which are independently consequential for individuals’ labor market outcomes, and therefore of independent scientific interest.
@@
The sequential nature of educational transitions implies that a causal mediation framework can be employed to study the causal paths by which education’s “value-added” occurs. Specifically, we can consider the first transition in a sequence of educational levels of interest as a treatment variable, $A$, and subsequent transitions as mediators that “transmit” the effects of the treatment and of prior transitions, $M_{k}$ ($1\leq k\leq K$). For example, if we are interested in the total effect of high school completion on earnings, we may ask to what extent this total effect operates indirectly, through the effects of college attendance and college completion (putative mediators) on earnings, or directly, through alternative causal pathways. The insight that the total causal effect of education can be decomposed into its direct and indirect effects opens up a range of important research and policy-oriented questions. For example, tracing to what extent an early-stage educational intervention boosts outcomes such as earnings via its promotion of subsequent educational attainment (its indirect effects), or via earnings directly, would enable policy-makers to discern what drives the intervention’s value and to hone subsequent policy (e.g. Hurwitz and Howell, 2014; Sullivan et al., 2019; Castleman et al., 2020; Bird et al., 2021; Dynarski et al., 2021; Black et al., 2023; Turner and Gurantz, 2024). Relatedly, if the early intervention’s effects are heterogeneous across demographic groups, assessing the intervention’s direct and indirect effects could guide researchers to aspects of the educational experience that either promote or inhibit upward mobility. Nevertheless, prior empirical approaches are not well-suited to answering these questions: a “years-of-schooling” approach captures the direct effect of each additional year of schooling, while the dichotomous approach conflates the direct and indirect effects. [^2] A further strand of literature, especially prominent in labor economics, explores labor market returns to horizontal aspects of differentiation within a given educational level (e.g. college selectivity, as well as specific colleges) or college types (e.g. Cohodes and Goodman, 2014; Goodman et al., 2017; Mountjoy and Hickman, 2021; Chetty et al., 2023; Eller, 2023). While my proposed framework prioritizes the effects of different levels of education, I discuss in the conclusion how the framework could be extended to accommodate multivariate mediators.
@@ 2.1 A Single Intermediate Transition
Using potential outcomes notation, let $M(a)$ denote an individual’s potential value of the mediator if their treatment status were set to $a$, and let $Y(a,m)$ denote that individual’s potential outcome if their treatment and mediator statuses were set to $a$ and $m$, respectively. With our set of potential outcomes ($\{Y(1),Y(0),Y(1,0),Y(1,1),Y(0,1)\}$), we can then define potential outcomes that involve “natural” values of the mediator. For instance, $Y(1,M(0))$ represents the earnings an individual would have if they completed high school ($A=1$) but their college attendance were fixed to the level it would have been had they not completed high school ($A=0$).
With this notation, we can define the “natural” decomposition of the average treatment effect as follows (Pearl, 2001; Imai et al., 2010):
@@
$=\text{NDE}(a)+\text{NIE}(1-a),$ (1)
for $a=0,1$, where $\text{NDE}(a)=\mathbb{E}[Y(1,M(a))-Y(0,M(a))]$, and $\text{NIE}(a)=\mathbb{E}[Y(a,M(1))-Y(a,M(0))]$. This decomposition allows a researcher to determine how much of the overall effect of high school completion is due to facilitating access to college, versus directly, net of college attendance.
Potential outcomes such as $Y(1,M(0))$ are central to mediation, but they create challenges in estimation and interpretation due to their “cross-world” nature. In particular, a unit cannot simultaneously complete high school and not complete high school at the same time, so $Y(1,M(0))$ is not observable, even in principle. This creates challenges for identification. In particular, it requires a “cross-world independence” assumption in order to identify the natural effect decomposition of the ATE: $Y(a^{\prime},m)\perp\!\!\!\perp M(a)\mid A=a,X$.
Many authors have expressed skepticism about this assumption since it requires that there are no (measured or unmeasured) post-treatment confounders (confounders that are affected by the treatment and which affect the mediator and outcome). An alternative approach to mediation analysis involves estimating a quantity known as the controlled direct effect (CDE), which captures the causal effect of a treatment when the mediator is fixed to a given value:
$\text{CDE}(m)=\mathbb{E}[Y(1,m)-Y(0,m)],$
for $m=0,1$. This quantity is attractive because it is identified under a weaker assumption than that required for quantities like $Y(1,M(0))$, one that allows for the existence of observed post-treatment confounders. A drawback of this approach is that it does not quantify the extent of mediation through $M$ and only enables a researcher to rule out the existence of alternative mediators other than $M$ (see Acharya et al. (2016)).
In the context of educational effects, it is possible to relax the usual cross-world assumption by leveraging a key structural feature of educational transitions. As I formalize in the following section, I assume that educational transitions are characterized by monotonicity. Here, this means that individuals who do not complete high school cannot attend college, or $M(0)=0$. This sequential nature of educational transitions therefore implies the following restricted set potential outcomes: $\{Y(1),Y(0),Y(1,0),Y(1,1)\}$ (i.e., ruling out $Y(0,1)$. Further, since by the composition assumption $Y(a)=Y(a,M(a))$ (VanderWeele and Vansteelandt, 2009), under monotonicity $Y(0)=Y(0,M(0))=Y(0,0)$.
We can then apply these restrictions to Equation 1 as follows (see Zhou, 2022a):
ATE $=\mathbb{E}[Y(1)-Y(0)]$
$=\underbrace{\mathbb{E}[Y(1,M(0))-Y(0,M(0)]}_{=\text{NDE}(0)}+\underbrace{\mathbb{E}[Y(1,M(1))-Y(1,M(0))]}_{=\text{NIE}(1)}$
$=\mathbb{E}[Y(1,0)-Y(0,0)]+\mathbb{E}[Y(1,M(1))-Y(1,0)]$
$=\mathbb{E}[Y(1,0)-Y(0,0)]+\mathbb{E}\big[M(1)[Y(1,1)-Y(1,0)]\big]$
@@
$=\underbrace{\Delta_{0}}_{A\to Y}+\underbrace{\pi_{\textup{1}}\Delta_{\textup{1}}+\eta_{\textup{1}}}_{A\to M\to Y},$ (3)
@@
where the third equality follows by monotonicity, the fourth, because $Y(1,M(1)-Y(1,0)=M(1)[Y(1,1)-Y(1,0)]+[1-M(1)]\cdot[Y(1,0)-Y(1,0)]=M(1)[Y(1,1-Y(1,0)]$, and the fourth, by rules of covariance. Here, $\Delta_{\textup{0}}$ and $\Delta_{1}$ denote the direct effects of the first and intermediate transitions on the outcome, $A\rightarrow Y$ and $M\rightarrow Y$, respectively, $\pi_{\textup{1}}$ denotes the total effect of the first transition on the intermediate transition $A\to M$, and $\eta_{\textup{1}}$ denotes the covariance between the effect of the initial transition on completion of the second and the effect of the second transition on $Y$. Specifically, $\eta_{\textup{1}}$ is positive if those who would attend college given high school completion (i.e., $M(1)=1$) benefit more from college attendance in terms of their later earnings (i.e., have a larger Y(1,1)Y(1,0)Y(1,1)-Y(1,0)) than those who do not (i.e., $M(1)=0$), and negative if the opposite is true. Meanwhile, the composite term $(\pi_{\textup{1}}\Delta_{1}+\eta_{\textup{1}})$ captures the average indirect effect of the treatment via the intermediate transition ($A\rightarrow M\rightarrow Y$), comprising the sum of (i) the probability of college enrollment if an individual graduated high school, multiplied by the direct of college enrollment, and (ii) the covariance between college enrollment and its direct effect on earnings.
While motivated by the natural effect decomposition of the ATE, the monotonicity constraint on $M$ leads to several differences from Equation 1. First, whereas Equation 1 can be written for $a=0$ and $a=1$, Equation 3 is the algebraically unique natural effect decomposition of the ATE under monotonicity. This is because the natural effect decomposition for $a=0$ hinges on the counterfactual term $Y(1,M(0))=Y(1,0)$, whereas the natural effect decomposition for $a=1$ hinges on the counterfactual term $Y(0,M(1))$. Under monotonicity, only the first of these quantities is well-defined. Second, under mediator monotonicity, $\text{NDE}(0)=\text{CDE}(0)$. Therefore, while in general the CDE does not quantify the extent of mediation through $M$ (since the difference between the ATE and CDE indicates the portions of the total effect due to interaction without mediation, and due to mediation, both with and without interaction (VanderWeele, 2014)) in this special case the CDE completely characterizes the extent of mediation. [^3] An alternative way to see this is that, as VanderWeele (2014) shows, the individual-level natural direct effect can be written as (Y(1,0)Y(0,0)+M(0)(Y(1,1)Y(0,1)Y(1,0)+Y(0,0))(Y(1,0)-Y(0,0)+M(0)(Y(1,1)-Y(0,1)-Y(1,0)+Y(0,0)), but under monotonicity, $M(0)=0$ for all individuals, and so the individual-level natural direct effect reduces to the first term. Interestingly, under monotonicity there is no “interaction effect” between the treatment and mediator (since the mediator can only take a value of 1 when the treatment is activated), and so the NIE reduces to pure indirect effect (PIE) discussed in VanderWeele (2014). Hence, under monotonicity, ATE = CDE + PIE, and there are no interaction terms in the decomposition. Third, and relatedly, because $Y(1,M(0))=Y(1,0)$, Equation 3 does not depend on any cross-world counterfactuals, and therefore can be identified under weaker assumptions that Equation 1.
Finally, the decomposition in Equation 3 can be compared with a “randomized effect” decomposition of treatment effects. An alternative approach to avoiding cross-world counterfactuals and treatment-induced confounding is to use randomized intervention analogues to the natural direct and indirect effects (rNDE and rNIE) (VanderWeele and Vansteelandt, 2014). Rather than setting the mediator to the level it would “naturally” take under an alternative treatment level, these estimands conceptualize interventions that randomly draw the mediator from its population distribution under each treatment regime.
Analogous to the natural effect decomposition in Equation 1, we can define a randomized effect decomposition of a treatment effect as follows:
rATE $=\text{rNDE}(a)+\text{rNIE}(1-a),$ (4)
for $a=0,1$, where $\text{rNDE}(a)=\mathbb{E}[Y(1,G_{a\mid X})-Y(0,G_{a\mid X})]$ and $\text{rNIE}(a)=\mathbb{E}[Y(a,G_{1\mid X})-Y(a,G_{0\mid X})]$, and rATE is defined to be the sum of these two components. Here, $G_{a\mid X}\equiv\text{Pr}(M=m|A=a,x)$ denotes a value of the mediator randomly drawn from its conditional distribution under treatment $A=a$ given baseline covariates $X$. Importantly, Equation 4 decomposes not the ATE but a randomized average treatment effect (rATE), which may differ from the ATE. By doing so, it avoids the cross-world counterfactual $Y(a,M(a^{\prime}))$ by substituting a randomized intervention on the mediator distribution. This substitution identifies the indirect and direct pathways, even when post-treatment confounders of the mediator–outcome relationship are present.
When $a=0$, the rNDE captures exactly the NDE under a monotonicity constraint (i.e., $\text{rNDE}(0)=\Delta_{0}$). This is because $\text{Pr}(M=1|A=0,x)=0$ and $\text{Pr}(M=0|A=0,x)=1$ and so the randomized effect reduces to the CDE. Using results from Yu et al. (2024), the difference between the NIE under monotonicity and the rNIE is equal to
$\pi_{1}\Delta_{1}+\eta_{1}-\text{rNIE}$ $=\mathbb{E}[\text{cov}[M(1)-M(0),Y(1,1)-Y(1,0)|X]]$
$=\mathbb{E}[\text{cov}[M(1),Y(1,1)-Y(1,0)|X]]\text{ (since }M(0)=0).$
This expression highlights that the NIE–rNIE gap measures residual within-$X$-stratum covariance between the mediator and its causal effect on $Y$, i.e., dependence induced by intermediate confounding. Thus, deviations between the ATE and rATE will stem from this within-stratum covariance. The randomized intervention framework parallels the monotonic decomposition in Equation 3, which eliminates logically impossible counterfactuals (e.g., $Y(0,M(1))$) through structural constraints on the mediator rather than redefining the estimand itself.
@@ 2.2 Generalization to $K$ Intermediate Transitions
I now generalize the approach introduced in the preceding section to the case of $K$ intermediate transitions. As previously, I denote the treatment (“initial transition”) of high school graduation by $A$, and use $M_{1},\dots M_{K}$ to refer to the $K$ subsequent transitions of interest (“intermediate transitions”), where I assume that all of $M_{1},\dots,M_{K}$ are binary and that for any $i<j$, $M_{i}$ temporally precedes $M_{j}$. For instance, we may wish to decompose the total effect of high school completion on earnings via college attendance ($M_{1}$), college completion ($M_{2}$) and graduate school attendance ($M_{3}$). Let an overbar denote a vector of variables, such that $\overline{M}_{k}=(M_{1},M_{2},\dots M_{k})$ and $(1,\overline{1}_{k-1})=(A=1,M_{1}=1,\dots,M_{k-1}=1)$. Further, let $[K]$ denote the set $\{0,1,\dots,K\}$. In addition, I denote by $X$ a vector of pretreatment confounders of the effect of $(A,\overline{M}_{k})$ on ($M_{k+1},Y)$, and by $\overline{Z}_{k}=(Z_{1},\dots Z_{k})$ a vector of intermediate confounders that may confound the causal effect of $M_{k}$ on ($M_{k+1},Y)$. Using potential outcomes notation, $Y(1,\overline{1}_{k-1},m_{k})$ thus denotes an individual’s potential earnings if they completed, possibly contrary to fact, the treatment and the first k1k-1 intermediate transitions, and then either completed ($m_{k}=1)$ or did not complete ($m_{k}=0$) the $k$th intermediate transition. Similarly, $M_{k+1}(1,\overline{1}_{k})$ denotes an individual’s potential value of the $k+1$th intermediate transition were that individual to complete the treatment and the first $k$ intermediate transitions. As is standard in the mediation literature, I make the following composition assumption (VanderWeele and Vansteelandt, 2009):
@@
Composition: $Y(1,\overline{1}_{k-1},m_{k})=Y(1,\overline{1}_{k-1},m_{k},M_{k+1}(1,\overline{1}_{k-1},m_{k})),\forall k\in[K-1],M_{0}\equiv A$.
In words, Assumption 1 states that a person’s potential outcome under $(1,\overline{1}_{k-1},m_{k})$ is equal to their potential outcome under $A=1,\dots,M_{k-1}=1,m_{k}$ and under the value $M_{k+1}$ would naturally take under $A=1,\dots,M_{k-1}=1,m_{k}$. I also invoke the following constraint on units’ potential transition values:
@@
Informally, Assumption 2 (monotonicity) states that an individual’s potential $k+1$th transition value is deterministically 0 if that individual fails to complete the prior ($k$th) transition. It is analogous to a one-sided non-compliance assumption within an instrumental variables (IV) framework, which precludes the presence of both “defiers" as well as “always-takers" principal strata. We can then use this assumption to decompose the ATE of $A$ on $Y$, which I denote by $\tau_{0}$. Specifically, let $\tau_{k}$ denote the gross effect of the $k$th mediator on $Y$, i.e.,
$\tau_{k}=\mathbb{E}[Y(1,\overline{1}_{k})-Y(1,\overline{1}_{k-1},0)],$
@@
$\Delta_{k}=\mathbb{E}[Y(1,\overline{1}_{k-1},1,0)-Y(1,\overline{1}_{k-1},0)].$
To explicate my approach, note that the gross effect of the $k$th mediator, $\tau_{k}$, includes not only the direct effect $M_{k}\rightarrow Y$, net of subsequent educational transitions $\Delta_{k}$, but also the indirect effects of $M_{k}$ via subsequent transitions ($M\rightsquigarrow Y$, where a squiggly arrow denotes a combination of multiple paths). This insight motivates us to further decompose $\tau_{k}$ into its direct and indirect components. Under the composition assumption, $\tau_{k}$ can be decomposed as
@@
$\pi_{k+1}$ $=\mathbb{E}[M_{k+1}(1,\overline{1}_{k})],$
$\eta_{k+1}$ $=\textup{cov}[M_{k+1}(1,\overline{1}_{k}),Y(1,\overline{1}_{k+1})-Y(1,\overline{1}_{k},0)].$